# Gamma-ray burst collimated fireballs: long vs short bursts by angular momentum and mass ratio

### v2

@@ Abstract
In this study, we investigate the gravitational collapses of rotating stellar systems accounting for Gamma-Ray Burst jet progenitors. Based on the virial theorem of hadron collisional relaxations and Newtonian slow-rotating approximation, we analyze the conversion of gravitational binding energy into kinetic energy of hadrons, whose collisions produce photons and electron-positron pairs forming fireballs. Our qualitative analysis implies that rotation effects collimate fireballs with nontrivial angular momenta along the propagating direction, thus possibly making ultra-relativistic jets. Results reveal the possible trends that the fireball becomes more collimated and the jet angle decreases as the total angular momentum and mass ratio $J/M$ of the slow-rotating collapsing core increases. Discussing the extrapolation of these trends to fast-rotating collapsing systems, we speculate that the ratio $J/M$ should be a key quantity for differentiating long bursts (massive core collapses) from short bursts (binary coalescence). We derive the intrinsic correlations of collimated fireball quantities that should be imprinted on a large sample of observed GRB data as empirical correlations.
@@ 1 Introduction
Gamma-ray bursts (GRBs) are the most energetic and complex events in the Universe. Within mere seconds, their progenitors release high-energy photons of tremendous isotropic equivalent energy estimated ranging from $10^{49}$ to $10^{54}$ ergs, assuming progenitors originate from opaque spherical fireballs of large particle number and energy densities [1]. The total energy budget should be smaller, considering that GRBs are highly collimated events, with their energy confined within a narrow ultra-relativistic collimated outflow (jet) [2, 3]. The typical jet opening angle is usually a few degrees [4, 5]. For example, the jet opening angle of GRB 170817A is estimated to be around 20 degrees [6, 7]. The jet opening angle of the brightest GRB 221009A is estimated to be exceptionally narrow, approximately 0.6 degrees [8, 9].
@@
Significant advancements in our comprehension of GRBs dynamics and phenomena have been achieved through extensive research efforts [10, 11, 12, 13, 14, 15, 16]. Among others, the ultra-relativistic hydrodynamical outflow of spherical fireballs has been studied well numerically and analytically, e.g., Refs. [10, 15, 17, 18, 19, 20, 21].
@@
However, what are the dominant dynamics of GRB progenitors that ultimately yield the formation of an energetic and optically thick plasma composed of photons and electron-positron pairs, dubbed as a collimated fireball, making an ultra-relativistic jet? There are numerous theoretical investigations in the direction following the Penrose process [22, 23] of extracting rotational energy from particles falling into a spinning black hole (SBH), and Blandford and Znajek’s subsequent work [24, 25] that jet energy could be extracted from the SBH rotational energy and transported out along large-scale magnetic fields surrounding SBH, making a relativistic jet.
@@
In contrast and comparison, we study the possibility of how significant amounts of stellar gravitational binding energy and angular momentum release to collimated fireballs in the gravitational collapse and coalescence of massive and rotating stellar systems. The critical questions are within just a few seconds during gravitational collapses: (i) how this immense gravitational energy is converted (extracted) into photon and pair energies; (ii) why photon and pair energies form opaque and collimated fireballs of huge energy and number densities; (iii) how collimated fireballs acquire nontrivial angular momentum and make ultra-relativistic jets. In Ref. [26], we study how the spherical fireball forms during gravitational homologous collapses of massive stellar cores. Here, considering the collapses of massive and rotating stellar systems, we study the formation of a collimated fireball with nontrivial angular momentum, whose hydrodynamical outflow makes an ultra-relativistic jet.
@@
In a self-gravitating system, Lynden-Bell and Wood [27, 28] pioneered the study of gravo-thermal catastrophe using the framework of violent hadron collisionless relaxation and the virial theorem. Their work demonstrated the relaxation process in which hadrons gain their mean kinetic heat energy from gravitational potential energy. Based on the local virial theorem and the homologous core collapse scenario [29], we show that hadrons gain mean kinetic heat energy from gravitational binding energy through collisional relaxation. On the other hand, the production of photons by hadron collisions has been extensively studied in the context of heavy-ion collisions [30, 31, 32, 33, 34, 35]. These two aspects imply that hadron-hadron collisions subsequently produce photons and pairs, converting their mean kinetic energy to photons’ energy. As a result, the gravitational binding energy converts into the energies of opaque photon-pair plasma (spherical fireball) in gravitational collapses [26]. This article investigates the possibility of forming collimated fireballs and ultra-relativistic jets in gravitational homologous collapses of massive rotating core collapses and binary coalescence.
@@
We organize this article as follows. Based on the adiabatic assumption, Sections 2, 3, and 4 present discussions on gravo-thermal catastrophe, the virial theorem, and photon production from hadron relaxations and collisions. In Section 5, we discuss the gravitational homologous collapse of a slowly rotating core. Sections 6 and 7 study collimated fireball and ultra-relativistic jet properties. Section 8 discusses the intrinsic correlations of collimated fireball properties and how they relate to the empirical correlations from GRB observations. Section 9 discusses the trends of collimated fireball properties as varying progenitor angular momentum and mass ratio, in connection with the characteristics of long versus short GRBs.
@@ 2 Adiabatic approximation in gravitational collapse
3 Local virial theorem and gravo-thermal configuration
We consider a stellar core composed of $N$ baryons of mass $m$ [^1] By the term baryons, we indicate nucleons, nuclei, hadrons, quark-gluon plasma that carry baryon numbers.. In gravitational collapse, baryons gain their mean kinetic energy $F$ from the gravitational binding energy $U$ via their collisionless relaxation in time-varying gravitational potential. This gravo-thermal catastrophe phenomenon of stellar cores is studied by considering violent relaxation [27] and equipartition theorem [28]. In these studies, authors considered the stellar core to be an isothermal core of baryon temperature $T\ll m$ and internal “heat” energy $F=\frac{3}{2}TN\ll mN$. It is in an equilibrium or equipartition configuration, where the Clausius virial theorem applies,
@@ 3 Local virial theorem and gravo-thermal configuration
These studies indicate that the baryon collisionless relaxation process in gravitational potential and under external pressure leads to two key consequences:
@@
The second point holds, provided the time scales of microscopic relaxation processes are much smaller than those of macroscopic gravitational and hydrodynamical processes, justifying adiabatic approximation. Following the discussions in the pioneer works [36, 37], it is noted that the gravitational evolution time scale due to stellar evaporation and consumption is longer than the core relaxation time scale, the core will maintain an approximate isothermal core profile (3.1) and evolve homologously.
@@ 4 Photon and $e^{+}e^{-}$ pair production from hadron collisions
The gravo-thermal catastrophe produces not only heat energy for massive baryons, given by $(3/2)T(\rho/m)dV$ (as in equation (3.3)), but also thermal energy for relativistic particles, described by $\rho_{\gamma}dV$ (as in equation (4.2)). For a baryon temperature $T$ on the order of ${\mathcal{O}}(10^{2})$ MeV, as will be shown later, the photon number density $n_{\gamma}$ (4.1) and energy density $\rho_{\gamma}$ (4.2) are very large. It implies that photons become opaque and thermalized by interacting with pairs of electrons and positrons. The local mean temperature $T_{\gamma}$ of thermalized photons can be estimated by equating photon thermal energy density $T^{4}_{\gamma}$ to the energy density $\rho_{\gamma}$ given in equation (4.2)
@@
This indicates that via processes $\gamma+\gamma\leftrightarrow e^{+}+e^{-}$, photons and electron-positron pairs participate in thermalisation or thermal equipartition in particle energy and number. As a result, these processes lead to an electrically neutral, deeply opaque, and thermalised plasma (fluid) of photons and electron-positron pairs, i.e., a “collimated fireball”. More detailed discussions can be found in Ref. [26].
@@ 5.1 Euler and Poisson equations in Newtonian approximation
We consider a slowly rotating stellar core undergoing a gravitational collapse process, whose rotation energy is much smaller than gravitational binding energy. Following the semi-analytical approach [29, 26], we will study the homologous gravitational collapse of slowly rotating and massive spherical cores by using the gravitational and centrifugal potentials in Newtonian approximation. We aim to explain how hadron collisions in gravitational potential balanced by centrifugal potential convert gravitational binding energy to the thermal energy of photons and pairs, creating a collimated fireball during gravitational collapses. Microscopic physical processes, e.g., photon production by hadron collisions, at a small spacetime scale are not affected much by using the general relativity (GR) or Newtonian framework at a macroscopic spacetime scale. However, it affects the spacetime distributions of collimated fireball properties, e.g., energy density, particularly approaching horizons. In this study, we attempt to obtain the approximate results and qualitative estimations of the total collimated fireball thermal energy, size and other quantities averaged over the entire collimated fireball volume and formation time at the end of collapses. We also consider that such approximate results should indicate a lower bound of counterpart energetic quantities obtained in the GR framework since the GR gravitational binding energy (potential) is larger than the Newtonian one in the gravitational collapse process from infinity to horizon $2M$. The numerical approach based on the GR framework is necessary for quantitative studies of collimated fireball formation from gravitational collapses of massive rotating stellar and binary systems. It is fundamental for the fireball’s inner part dynamics near the horizon. Nevertheless, due to the complex physical issues of GRB progenitors, we first figure out what macroscopic and microscopic processes are dominant and crucial for producing and collimating fireballs. Therefore, it deserves preliminary analysis using the analytical approach based on the Newtonian approximation in the slow-rotating limit.
In the Newtonian approximation, we describe the process by the continuity equation for the baryon number conservation, Euler’s equation for energy-momentum conservation, Poisson’s equation for gravitational potential $\phi$ are,
$\frac{\partial\rho}{\partial t}+{\bf\nabla}\cdot(\rho\bf{u})$ $=$ $0,$ (5.1)
$\Big{(}\frac{\partial{\bf u}}{\partial t}\Big{)}+{\bf\nabla}\Big{(}\frac{1}{2}|{\bf u}|^{2}\Big{)}+(\nabla\times{\bf u})\times{\bf u}+\nabla h+{\bf\nabla}\phi+{\bf\nabla}\phi_{c}$ $=$ $0,$ (5.2)
$\nabla^{2}\phi-4\pi G\rho$ $=$ $0,$ (5.3)
@@
We formally add the gradient of centrifugal potential $\phi_{c}$ given by ${\bf\nabla}\phi_{c}=j{\bf\nabla}\Omega$, where $j$ and $\Omega$ are angular momentum and velocity respectively. These equations reduce to their counterparts in the hydrodynamic equilibrium case, describing the core’s state when the dynamic effects of collapse are not present.
@@ 5.2 Static equilibrium configuration and rotation law
where $x=R\sin\theta$ and $z=R\cos\theta$ are the cylindrical coordinates, $R$ is the radial coordinate, and $\theta$ is the zenith (polar) angle. The rotating axis is along the $\hat{z}$ direction. The system is axially symmetric and we omit the coordinate $y$ in formulae. The centrifugal potential is given by
@@
which is uniform across the space. The constant on the right-handed side relates to the material binding energy. To generalize this to a model of a rigidly and slowly rotating stellar core undergoing homologous collapse, we need to consider the density, velocity, and potential changes during the collapse.
@@ 5.4 Spherical and homologous solution in space
where the positive eigenvalue $\lambda$ relates to the material binding energy of baryon fluid per energy density $\rho$ [29] or external pressure acting on the core surface. Namely, $\lambda dm$ relates to an elementary material binding energy for a baryon fluid element of mass $dm$ around the nuclear density. Equation (5.17) describes the time evolution of the length scale function $a(t)$ in the collapsing core. The time-independent homologous spatial solution to the Euler’s equation (5.16) is
@@
The mean thermal index $\gamma$ in the simple polytropic EoS (3.4), depending on the microscopic physics of material density and interactions, represents the average effect of repulsive heat energy pressure against attractive gravity force. Small $\gamma$ values for small heat pressure imply stellar cores would be unstable and undergo gravitationally collapses. Goldreich and Weber [29] show that such EoS (3.4) and $\gamma=4/3$ provide a reasonable approximation to collapsing stellar cores during the early phase before reaching nuclear density. In this simplified model, they obtain the homologous collapse solution of the onset value $\gamma=4/3$, below which massive cores undergo gravitational instability and collapse.
Following the same approach, we obtain [26] homologous collapse solutions for general values $31<\gamma<4/$ 3, studying how spherical fireballs form in gravitational collapses. However, the realistic situations of gravitational collapses of massive cores approaching nuclear density are much more complex due to nucleon interactions and possible phase transitions at different densities. These are beyond the scope of present studies. In this article, we approximately consider EoS (3.4) and the mean thermal index $\gamma\gtrsim 1$ as an effective $\mathcal{O}(1)$ free parameter relevant to how much the gravitational binding energy gains for forming fireballs. The larger (smaller) $\gamma$ value is, and the smaller (larger) gravitational binding energy gains. We select three $\gamma$ values (5.20) for illustration.
@@
orange $1.24$ $1.0\times 10^{-4}$ $23$
blue $1.23$ $8.0\times 10^{-5}$ $36$
green $1.225$ $8.0\times 10^{-6}$ $34$
@@
Figure 1: Colours orange, blue and green on lines for three different eigenvalues selected in the list (5.20). We show numerical solutions to the eigenvalue problem of Eqs. (5.16) and (5.17). The homologous density profile $f^{3}(r)$ (5.12) is independent of the time. We plot it as a function of dimensionless radius $r=R/a$. It vanishes $f^{3}(r_{s})=0$ at the core outer radius $r_{s}$, see values in the list (5.20). The increase of the core density $\rho=\rho_{c}(t)f^{3}$ (5.12) in homologous collapse is due to the increase of the centre density $\rho_{c}(t)$ via the decrease of the length scale function $a(t)$ (5.14) and Fig. 2.

### v3

# Collimated and spinning fireballs for ultra-relativistic jets: long vs short Gamma-ray bursts by angular momentum and mass ratio

@@ Abstract
In this study, we investigate the gravitational collapses of rotating stellar systems accounting for Gamma-Ray Burst jet progenitors. Based on the virial theorem of hadron collisional relaxations and Newtonian slow-rotating approximation, we analyze the conversion of gravitational binding energy into kinetic energy of hadrons, whose collisions produce photons and electron-positron pairs forming fireballs. Our qualitative analysis implies that rotation effects collimated and spinning fireballs with nontrivial angular momenta along the propagating direction, thus making ultra-relativistic jets. Results reveal the possible trends that the fireball becomes more collimated and the jet angle decreases as the total angular momentum and mass ratio $J/M$ of the slow-rotating collapsing core increases. Discussing the extrapolation of these trends to fast-rotating collapsing systems, we speculate that the ratio $J/M$ should be a key quantity for differentiating long bursts (massive core collapses) from short bursts (binary coalescence). We derive the intrinsic correlations of collimated fireball quantities that should be imprinted on a large sample of observed GRB data as empirical correlations.
@@ 1 Introduction
Gamma-ray bursts (GRBs) are the most energetic and complex events in the Universe. Within mere seconds, their progenitors release high-energy photons of tremendous isotropic equivalent energy estimated to range from $10^{49}$ to $10^{54}$ ergs, assuming progenitors originate from opaque spherical fireballs of large particle number and energy densities [1]. The total energy budget should be smaller, considering that GRBs are highly collimated events, with their energy confined within a narrow ultra-relativistic collimated outflow (jet) [2, 3]. The typical jet opening angle is usually a few degrees [4, 5]. For example, the jet opening angle of GRB 170817A is estimated to be around 20 degrees [6, 7]. The jet opening angle of the brightest GRB 221009A is estimated to be exceptionally narrow, approximately 0.6 degrees [8, 9].
@@
Significant advancements in our comprehension of GRBs dynamics and phenomena have been achieved through extensive research efforts [10, 11, 12, 13, 14, 15, 16]. Among others, the ultra-relativistic hydrodynamical outflow of spherical fireballs has been studied well numerically and analytically, e.g., Refs. [10, 15, 17, 18, 19, 20, 21]. Recently, numerical investigations have made notable progress in understanding jet evolution from early launching to later propagation and interaction with ambient stellar medium away from jet progenitors, e.g., Refs. [22, 23, 24, 25, 26, 27, 28, 29, 30, 31].
@@
However, what are the dominant dynamics of GRB progenitors that ultimately yield the formation of an energetic and optically thick plasma composed of photons and electron-positron pairs, dubbed a collimated and spinning fireball, making an ultra-relativistic jet? There are numerous theoretical investigations in the direction following the Penrose process [32, 33] of extracting rotational energy from particles falling into a spinning black hole (SBH), and Blandford and Znajek’s subsequent work [34, 35] that jet energy could be extracted from the SBH rotational energy and transported out along large-scale magnetic fields surrounding SBH, making a relativistic jet.
@@
In contrast and comparison, we study the possibility of how significant amounts of stellar gravitational binding energy and angular momentum are released to collimated and spinning fireballs in the gravitational collapse and coalescence of massive and rotating stellar systems. The critical questions are within just a few seconds during gravitational collapses: (i) how this immense gravitational energy is converted (extracted) into photon and pair energies; (ii) why photon and pair energies form opaque and collimated fireballs of huge energy and number densities; (iii) how collimated fireballs acquire nontrivial angular momentum and make ultra-relativistic jets. In Ref. [36], we study how the spherical fireball forms during gravitational homologous collapses of massive stellar cores. Here, considering the collapses of massive and rotating stellar systems, we study the formation of a collimated fireball with nontrivial angular momentum, whose hydrodynamical outflow makes an ultra-relativistic jet.
@@
In a self-gravitating system, Lynden-Bell and Wood [37, 38] pioneered the study of gravothermal catastrophe using the framework of violent hadron collisionless relaxation and the virial theorem. Their work demonstrated the relaxation process in which hadrons gain their mean kinetic heat energy from gravitational potential energy. Based on the local virial theorem and the homologous core collapse scenario [39], we show that hadrons gain mean kinetic heat energy from gravitational binding energy through collisional relaxation. On the other hand, the production of photons by hadron collisions has been extensively studied in the context of heavy-ion collisions [40, 41, 42, 43, 44, 45]. These two aspects imply that hadron-hadron collisions subsequently produce photons and pairs, converting their mean kinetic energy to photons’ energy. As a result, the gravitational binding energy converts into the energies of opaque photon-pair plasma (spherical fireball) in gravitational collapses [36]. This article investigates the possibility of forming collimated fireballs and ultra-relativistic jets in gravitational homologous collapses of massive rotating core collapses and binary coalescence.
@@
We organize this article as follows. Based on the adiabatic assumption, Sections 2, 3, and 4 present discussions on gravothermal catastrophe, the virial theorem, and photon production from hadron relaxations and collisions. In Section 5, we discuss the gravitational homologous collapse of a slowly rotating core. Sections 6 and 7 study collimated fireball and ultra-relativistic jet properties. Section 8 discusses the intrinsic correlations of collimated fireball properties and how they relate to the empirical correlations from GRB observations. Section 9 discusses the trends of collimated fireball properties as varying progenitor angular momentum and mass ratio, in connection with the characteristics of long versus short GRBs.
@@ 2 Adiabatic approximation in gravitational collapse
3 Local virial theorem and gravothermal configuration
We consider a stellar core composed of $N$ baryons of mass $m$ [^1] By the term baryons, we indicate nucleons, nuclei, hadrons, quark-gluon plasma that carry baryon numbers.. In gravitational collapse, baryons gain their mean kinetic energy $F$ from the gravitational binding energy $U$ via their collisionless relaxation in time-varying gravitational potential. This gravothermal catastrophe phenomenon of stellar cores is studied by considering violent relaxation [37] and equipartition theorem [38]. In these studies, authors considered the stellar core to be an isothermal core of baryon temperature $T\ll m$ and internal “heat” energy $F=\frac{3}{2}TN\ll mN$. It is in an equilibrium or equipartition configuration, where the Clausius virial theorem applies,
@@ 3 Local virial theorem and gravothermal configuration
These studies indicate that the baryon collisionless relaxation process in the gravitational potential and under external pressure leads to two key consequences:
@@
The second point holds, provided that the time scales of microscopic relaxation processes are much smaller than those of macroscopic gravitational and hydrodynamical processes, justifying the adiabatic approximation. Following the discussions in the pioneer works [46, 47], it is noted that the gravitational evolution time scale due to stellar evaporation and consumption is longer than the core relaxation time scale, the core will maintain an approximate isothermal core profile (3.1) and evolve homologously.
@@ 4 Photon and $e^{+}e^{-}$ pair production from hadron collisions
The gravothermal catastrophe produces not only heat energy for massive baryons, given by $(3/2)T(\rho/m)dV$ (as in equation (3.3)), but also thermal energy for relativistic particles, described by $\rho_{\gamma}dV$ (as in equation (4.2)). For a baryon temperature $T$ on the order of ${\mathcal{O}}(10^{2})$ MeV, as will be shown later, the photon number density $n_{\gamma}$ (4.1) and energy density $\rho_{\gamma}$ (4.2) are very large. It implies that photons become opaque and thermalized by interacting with pairs of electrons and positrons. The local mean temperature $T_{\gamma}$ of thermalized photons can be estimated by equating photon thermal energy density $T^{4}_{\gamma}$ to the energy density $\rho_{\gamma}$ given in equation (4.2)
@@
This indicates that via processes $\gamma+\gamma\leftrightarrow e^{+}+e^{-}$, photons and electron-positron pairs participate in thermalisation or thermal equipartition in particle energy and number. As a result, these processes lead to an electrically neutral, deeply opaque, and thermalized plasma (fluid) of photons and electron-positron pairs, i.e., a “collimated fireball”. More detailed discussions can be found in Ref. [36].
@@
We add a few remarks to end this section. The photon number and energy densities (4.1) and (4.2) and efficiency of producing such densities should be underestimated (lower bounds) in hadron collisions at densities $(\rho\sim 10\rho_{0})$. From the timescales of high-energy photons’ interacting dynamics and kinematics, we expect the exponential multiplicity of photon productions, like QED cascade [48, 49], to occur via photon-photon and photon-electron collisions, as well as photons, charged particles and magnetic field (virtual photons) interactions. At higher densities $n\gtrsim 10n_{0}$, a phase transition from hadrons (baryons) to quark-gluon plasma should occur. In this case, the baryon temperature $T$ (3.5) should be the temperature of the quark-gluon plasma. We expect even more efficient photon production multiplicity via QCD cascade processes, see Refs. [50, 51], to occur in reality. In these high-energy photons’ cascade processes, the total photon number $N_{\gamma}$ should be proportional to an exponential factor $e^{t^{\rm mac}_{\gamma}/t^{\rm mic}_{\gamma}}\gg 1$ and $t^{\rm mac}_{\gamma}>t^{\rm mic}_{\gamma}$, where $t^{\rm mic}_{\gamma}$ is the high-energy photon interacting timescale for particle productions and $t^{\rm mac}_{\gamma}$ is the timescale relating to the relaxation time (3.2) of hadron collisions in gravitational collapses. These topics deserve further investigation.
@@ 5.1 Euler and Poisson equations in Newtonian approximation
We consider a slowly rotating stellar core undergoing a gravitational collapse process, whose rotation energy is much smaller than the gravitational binding energy. Following the semi-analytical approach [39, 36], we will study the homologous gravitational collapse of slowly rotating and massive spherical cores by using the gravitational and centrifugal potentials in Newtonian approximation. We aim to explain how hadron collisions in gravitational potential, balanced by centrifugal potential, convert gravitational binding energy to the thermal energy of photons and pairs, creating a collimated fireball during gravitational collapses. Microscopic physical processes, e.g., photon production by hadron collisions, at a small spacetime scale, are not affected much by using the general relativity (GR) or Newtonian framework at a macroscopic spacetime scale. However, it affects the spacetime distributions of collimated fireball properties, e.g., energy density, particularly approaching horizons. In this study, we attempt to obtain the approximate results and qualitative estimations of the total collimated fireball thermal energy, size and other quantities averaged over the entire collimated fireball volume and formation time at the end of collapse. We also consider that such approximate results should indicate a lower bound of counterpart energetic quantities obtained in the GR framework, since the GR gravitational binding energy (potential) is larger than the Newtonian one in the gravitational collapse process from infinity to the horizon $2M$. The numerical approach based on the GR framework is necessary for quantitative studies of collimated fireball formation from gravitational collapses of massive rotating stellar and binary systems. It is fundamental for the fireball’s inner part dynamics near the horizon. Nevertheless, due to the complex physical issues of GRB progenitors, we first figure out what macroscopic and microscopic processes are dominant and crucial for producing and collimating fireballs. Therefore, it deserves preliminary analysis using the analytical approach based on the Newtonian approximation in the slow-rotating limit. The GR-based numerical simulation should not qualitatively change the general properties of physical effects and trends obtained in the Newtonian approximation.
In the Newtonian approximation, we describe the process by the continuity equation for the baryon number conservation, Euler’s equation for energy-momentum conservation, and Poisson’s equation for gravitational potential $\phi$ are,
$\frac{\partial\rho}{\partial t}+{\bf\nabla}\cdot(\rho\bf{u})$ $=$ 0 0,0 , (5.1)
$\Big{(}\frac{\partial{\bf u}}{\partial t}\Big{)}+{\bf\nabla}\Big{(}\frac{1}{2}|{\bf u}|^{2}\Big{)}+(\nabla\times{\bf u})\times{\bf u}+\nabla h+{\bf\nabla}\phi+{\bf\nabla}\phi_{c}$ $=$ 0 0,0 , (5.2)
$\nabla^{2}\phi-4\pi G\rho$ $=$ 0 0,0 , (5.3)
@@
We formally add the gradient of centrifugal potential $\phi_{c}$ given by ${\bf\nabla}\phi_{c}=j{\bf\nabla}\Omega$, where $j$ and $\Omega$ are angular momentum and velocity, respectively. These equations reduce to their counterparts in the hydrodynamic equilibrium case, describing the core’s state when the dynamic effects of collapse are not present.
@@ 5.2 Static equilibrium configuration and rotation law
where $x=R\sin\theta$ and $z=R\cos\theta$ are the cylindrical coordinates, $R$ is the radial coordinate, and $\theta$ is the zenith (polar) angle. The rotating axis is along the $\hat{z}$ direction. The system is axially symmetric, and we omit the coordinate $y$ in formulae. The centrifugal potential is given by
@@
which is uniform across the space. The constant on the right-hand side relates to the material binding energy. To generalize this to a model of a rigidly and slowly rotating stellar core undergoing homologous collapse, we need to consider the density, velocity, and potential changes during the collapse.
@@ 5.4 Spherical and homologous solution in space
where the positive eigenvalue $\lambda$ relates to the material binding energy of the baryon fluid per energy density $\rho$ [39] or external pressure acting on the core surface. Namely, $\lambda dm$ relates to an elementary material binding energy for a baryon fluid element of mass $dm$ around the nuclear density. Equation (5.17) describes the time evolution of the length scale function $a(t)$ in the collapsing core. The time-independent homologous spatial solution to Euler’s equation (5.16) is
@@
The mean thermal index $\gamma$ in the simple polytropic EoS (3.4), depending on the microscopic physics of material density and interactions, represents the average effect of repulsive heat energy pressure against the attractive gravity force. Small $\gamma$ values for small heat pressure imply that stellar cores would be unstable and undergo gravitational collapse. Goldreich and Weber [39] show that such EoS (3.4) and $\gamma=4/3$ provide a reasonable approximation to collapsing stellar cores during the early phase before reaching nuclear density. In this simplified model, they obtain the homologous collapse solution of the onset value $\gamma=4/3$, below which massive cores undergo gravitational instability and collapse.
Following the same approach, we obtain [36] homologous collapse solutions for general values $1<\gamma<4/3$, studying how spherical fireballs form in gravitational collapses. However, the realistic situations of gravitational collapses of massive cores approaching nuclear density are much more complex due to nucleon interactions and possible phase transitions at different densities. These are beyond the scope of the present study. In this article, we approximately consider EoS (3.4) and the mean thermal index $\gamma\gtrsim 1$ as an effective $\mathcal{O}(1)$ free parameter relevant to how much the gravitational binding energy gains for forming fireballs. The larger (smaller) $\gamma$ value is, the smaller (larger) gravitational binding energy gains. We select three $\gamma$ values (5.20) for illustration.
@@
orange 1.241.241.24 $1.0\times 10^{-4}$ 232323
blue 1.231.231.23 $8.0\times 10^{-5}$ 363636
green 1.2251.2251.225 $8.0\times 10^{-6}$ 343434
@@
Figure 1: Colours orange, blue and green on lines for three different eigenvalues selected in the list (5.20). We show numerical solutions to the eigenvalue problem of Eqs. (5.16) and (5.17). The homologous density profile $f^{3}(r)$ (5.12) is independent of time. We plot it as a function of dimensionless radius $r=R/a$. It vanishes $f^{3}(r_{s})=0$ at the core outer radius $r_{s}$, see values in the list (5.20). The increase of the core density $\rho=\rho_{c}(t)f^{3}$ (5.12) in homologous collapse is due to the increase of the centre density $\rho_{c}(t)$ via the decrease of the length scale function $a(t)$ (5.14) and Fig. 2.